# Transforming LLMs into Cross-modal and Cross-lingual Retrieval Systems

### v1

@@ 2 Method
We train a transformer-based DE model that encodes speech and text given a dataset, $\emph{D}=\{(x_{i},y_{i})\}$, where $x_{i}$ is a speech utterance and $y_{i}$ is its transcription. We denote the speech and text embeddings as $\bm{x_{i}}=E(x_{i})$ and $\bm{y_{i}}=E(y_{i})$, respectively, where $E$ is a transformer-based DE that encodes speech and text.
@@ 4 Model
Figure 1 shows an illustration of our model. We initialize our dual encoder from PaLM 2 XXS Google et al. (2023) and append a linear projection layer after pooling the outputs along the sequence length dimension. The embedding and linear projection layers are initialized randomly. After initializing our model from PaLM 2, we use a contrastive loss Hadsell et al. (2006). Appendix A.1 includes more details on our training setup.
@@ 5 Experiments
PaLM 2 XXS 86.15 13.85
mSLAM 76.9 14.6
Table 1: PaLM 2 results for $R@1$ and WER compared against the mSLAM on 102 languages from FLEURS for speech-to-text retrieval (S2T).
@@ 5.1 Speech-to-Text Retrieval
Table 1 shows the average $R@1$ and WER for S2T for 102 languages from FLEURS. We compare against the mSLAM model from Conneau et al. (2023), a model trained on 426k hours of S2T data in 51 languages and fine-tuned on FLEURS training data. Our model significantly outperforms the mSLAM baseline in $R@1$ and WER metrics despite being trained with only 1/10 of the data and having been initialized from a text-only LLM. More importantly, our model was only trained on the 21 languages in CoVoST-2 and never fine-tuned on the FLEURS training data.
@@ 5.1.2 Language Group Breakdown
Language Group (#) PaLM 2 (%) mSLAM (%)
Afro-Asiatic (7) 90.82 73.67
Atlantic-Congo (15) 79.47 86.77
Austro-Asiatic (3) 41.71 47.90
Austronesian (6) 85.74 75.50
Dravidian (4) 90.46 65.70
Indo-European (51) 92.38 84.62
Japonic (1) 63.23 5.80
Kartvelian (1) 74.57 70.50
Koreanic (1) 45.81 5.20
Kra-Dai (1) 37.12 3.20
Mongolic (1) 98.10 70.70
Nilo-Saharan (1) 94.53 91.00
Sino-Tibetan (3) 61.91 3.40
Turkic (4) 91.89 81.28
Uralic (3) 93.38 91.40
@@
Table 2 shows the $R@1$ language group breakdown for S2T on FLEURS. We find that although we only trained on 21 languages, our model significantly outperforms mSLAM in 13 of the 15 language groups. These results are consistent with the experiments in Hassid et al. (2023) which explore the effect of initializing speech language models from pre-trained LLMs.
@@ 5.2 Evaluating on Cross-Modal and Cross-Lingual Tasks
Figure 3: BLEU scores for FLEURS zero-shot S2TT when training on Transcripts or Transcripts + Translations. Combining transcripts and translation data improves zero-shot S2TT retrieval.
@@ 8 Acknowledgements
We would like to thank Shankar Kumar and Ankur Bapna for the valuable feedback on the draft of the paper. Chris Tar and Mario Guajardo-Céspedes for the early experiment discussions and feedback. Christian Frank, Duc Dung Nguyen, Alex Tudor, and Dalia El Badawy for helping answer questions about AudioPaLM.

### v2

# Using LLMs as Speech-to-Text Retrieval Systems

@@ 2 Method
We train a transformer-based DE model that encodes speech and text given a dataset $\emph{D}=\{(x_{i},y_{i})\}$, where $x_{i}$ is a speech utterance and $y_{i}$ is its transcription. We denote the speech and text embeddings as $\bm{x_{i}}=E(x_{i})$ and $\bm{y_{i}}=E(y_{i})$, respectively, where $E$ is a transformer-based DE that encodes speech and text.
@@ 4 Model
Figure 1 shows an illustration of our model. We initialize our dual encoder from PaLM 2 XXS Google et al. (2023) and append a linear projection layer after pooling the outputs along the sequence length dimension. The embedding and linear projection layers are initialized randomly. After initializing our model from PaLM 2, we use a contrastive loss Hadsell et al. (2006). Appendix A.1 includes more details on our training setup. We will refer to our proposed model as PaLM 2 DE.
@@ 5 Experiments
mSLAM DE Conneau et al. (2023) 76.9 14.6
PaLM 2 DE (Proposed Model) 86.15 13.85
Table 1: PaLM 2 DE results for R@1 and WER compared against the mSLAM DE on 102 languages from FLEURS for speech-to-text retrieval (S2T).
@@ 5.1 Speech-to-Text Retrieval
Table 1 shows the average $R@1$ and WER for S2T for 102 languages from FLEURS. We compare against the mSLAM DE model from Conneau et al. (2023), a model trained on 426k hours of S2T data in 51 languages and fine-tuned on FLEURS training data. Our model significantly outperforms the mSLAM DE baseline in $R@1$ and $WER$ metrics despite being trained with only 1/10 of the data and having been initialized from a text-only LLM. More importantly, our model was only trained on the 21 languages in CoVoST-2 and never fine-tuned on the FLEURS training data.
@@ 5.1.2 Language Group Breakdown
Language Group (#) mSLAM DE PaLM 2 DE
Conneau et al. (2023) (Proposed Model)
Afro-Asiatic (7) 73.67 90.82
Atlantic-Congo (15) 86.77 79.47
Austro-Asiatic (3) 47.90 41.71
Austronesian (6) 75.50 85.74
Dravidian (4) 65.70 90.46
Indo-European (51) 84.62 92.38
Japonic (1) 5.80 63.23
Kartvelian (1) 70.50 74.57
Koreanic (1) 5.20 45.81
Kra-Dai (1) 3.20 37.12
Mongolic (1) 70.70 98.10
Nilo-Saharan (1) 91.00 94.53
Sino-Tibetan (3) 3.40 61.91
Turkic (4) 81.28 91.89
Uralic (3) 91.40 93.38
@@
Table 2 shows the $R@1$ language group breakdown for S2T on FLEURS. We find that although we only trained on 21 languages, our model significantly outperforms mSLAM DE in 13 of the 15 language groups. These results are consistent with the experiments in Hassid et al. (2023) which explore the effect of initializing speech language models from pre-trained LLMs.
@@ 5.2 Evaluating on Cross-Modal and Cross-Lingual Tasks
Figure 3: BLEU scores for FLEURS zero-shot S2TT when training on Transcripts or Transcripts + Translations for PaLM 2 DE. Combining transcripts and translation data improves zero-shot S2TT retrieval.
@@ 8 Acknowledgements
We would like to thank Shankar Kumar and Ankur Bapna for the valuable feedback on the draft of the paper. Chris Tar, Mario Guajardo-Céspedes, and Jason Riesa for the early experiment discussions and feedback. Christian Frank, Duc Dung Nguyen, Alex Tudor, and Dalia El Badawy for helping answer questions about AudioPaLM.